# Article Type: Pre-Print

### v1

@@ Abstract
Cybersecurity threats in automotive systems pose significant risks to safety and reliability. This article introduces a methodology integrating threat-informed dynamic security modelling with a Threat Analysis and Risk Assessment workflow. Using the example of an In-Vehicle Infotainment system, we demonstrate the methodology’s application in risk management to strengthen automotive resiliency.
Keywords: risk management, security, autonomous vehicles
journal: IEEE Security & Privacy
Introduction
Automotive systems have transformed into sophisticated platforms brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer seamless external connectivity through Bluetooth, Wi-Fi, and GPS. Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong (2021) observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
@@ Introduction
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. In our previous work [1], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example. Our investigation yielded several findings, including:
@@
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security, overlooking critical aspects of how such breaches can directly affect automotive functions and safety. In other words, threat modelling of automotive systems often do not consider consequences of cybersecurity attacks beyond information security. On a similar note, attack models designed with traditional IT systems in mind do not fully align with multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions.
In short, our investigation has highlighted limitations in current threat and attack modelling methods that underscore the need for a tailored, dynamic approach. Threat and attack models need to iteratively and continuously consider both consequences and characteristics of cyber-physical attacks so as to offer a comprehensive approach to safeguarding automotive systems.
Contributions. In this article, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — a novel security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
Unlike existing security modelling approaches, ACTISM dynamically updates security models with new information acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ significantly from those in traditional IT systems. These incidents can lead to complex failure modes with both cyber and physical consequences, such as vehicle misnavigation and potentially dangerous outcomes. By adopting a consequence-driven and cyber-informed approach to cybersecurity, ACTISM ensures that cyber-physical attacks, effects and consequences are considered in security modelling for automotive systems.
ACTISM seamlessly and systematically integrates with, and adapts to a number of threat analysis and risk assessment frameworks such as:
@@
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method for automotive systems.
@@
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities.
@@
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks. Furthermore, we illustrate ACTISM’s adaptive response to a real-world vulnerability disclosure affecting the IVI system, underscoring its effectiveness in addressing emerging threats and showcasing its potential applicability to other automotive systems.
Finally, we outline avenues for future research and development in this area.
Related Work
Various articles have attempted to address security modelling in automotive systems.
@@
Abuabed et al. [3] proposed a framework for assessing the cybersecurity vulnerabilities of automotive vehicles. Their approach employed STRIDE to classify threats. For assessing the impact of threats, they used version 1.0 of the HEAVENS security model [4], and they evaluated attack feasibility with CVSS version 2.0. Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS). Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths, facilitating a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
The framework has certain limitations. Firstly, the threat identification process did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This may arise from the assumption that physical attacks do not involve data flows. However, physical attack vectors, such as "USB Man-in-the-Middle (USB MITM)" attacks, can exploit and facilitate data flows by intercepting and altering communication. As a result, neglecting physical attack vectors introduces a significant gap in the overall security design of the vehicle.
Secondly, both the HEAVENS version 1.0 and CVSS version 2.0 frameworks adopted in Abuabed et al. [3] are no longer current. Their newer versions include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [5] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. Additionally, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [6], who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
@@
Threat-informed Dynamic Security Modelling with ACTISM
@@
We propose ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example.
Unlike conventional security modelling methodologies, ACTISM seamlessly integrates both threat and attack modelling in a continuous, iterative workflow throughout all stages of the vehicle’s life cycle. Moreover, ACTISM is dynamic, receiving updates through cybersecurity threat information acquired from vulnerability advisories and ongoing security monitoring.
Figure 1: Overview of threat-informed dynamic security modelling
As illustrated in Figure 1, a threat-informed dynamic security modeling approach relies on the seamless integration of threat and attack modeling, utilising methodologies and tools like Microsoft’s STRIDE and CVSS 3.1. A key component of this approach is the strategic incorporation of threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
@@
Such an iterative modelling approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [8], breaching the trust boundary within security models.
@@
ACTISM does not replace threat and attack modelling. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By incorporating an iterative modelling approach, the methodology enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM enhances the TARA framework by dynamically updating the cybersecurity posture of automotive systems based on the evolving threat landscape and the specific impacts of potential attacks on system functionality and safety. ACTISM comprises eight key steps, outlined as follows:
  1. 1.
    Identify assets
  2. 2.
    Construct Data Flow Diagrams
  3. 3.
    Model threats
  4. 4.
    Evaluate impact
  5. 5.
    Model attacks
  6. 6.
    Evaluate attack feasibility
  7. 7.
    Evaluate risks
  8. 8.
    Review, update, and iterate
The details of each step will be elaborated on in the next section, using the running example of an IVI system.
Lastly, we further incorporate the following enhancements in ACTISM to increase its practical utility amongst cybersecurity practitioners:
  • Threat analysis: We employ the STRIDE methodology using a STRIDE-per-element approach where threats are assessed against each element in the Data Flow Diagram. We do not use a STRIDE-per-interaction approach, as this yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [9]. Additionally, we use HEAVENS version 2.0 [10] for impact assessment, as version 2.0 offers enhancements over version 1.0, providing more accurate and comprehensive risk assessments.
  • Attack analysis: We use CVSS version 3.1 to assess attack feasibility ratings, allowing us to leverage current cybersecurity vulnerability knowledge bases. Additionally, CVSS version 3.1 offers granular metrics and scoring options, which consider factors affecting the exploitability and impact of a vulnerability.
Applying ACTISM to an In-Vehicle Infotainment (IVI) System
@@
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing seamless connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
@@
Step 1: Identify assets
@@
In this article, we reference the IVI HPC system DFD from Das et al. (2024) [2] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling.
@@
Step 2: Construct Data Flow Diagrams
@@
Step 3: Model threats
@@
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Figure 3.
Figure 3: Threats generated by Microsoft Threat Modelling Tool
Step 4: Evaluate impact
To analyse the impact of the identified threats, we use the HEAVENS version 2.0 framework [10] which is specifically designed for automotive risk assessments. This framework allows us to systematically evaluate the potential impacts across four key dimensions: safety (s), financial (f), operational (o), and privacy and legislation (pl).

### v2

# ACTISM: Threat-Informed Dynamic Security Modelling for Automotive Systems

@@ Abstract
Cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety. This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge. We also outline avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
Keywords:
Risk management Security Automotive systems
1 Introduction
Automotive systems have transformed into complex cyber-physical systems brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer external connectivity through Bluetooth, Wi-Fi, and GPS. In addition, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong [11] observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
@@ 1 Introduction
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
  • Interconnectivity: Modern vehicles are highly interconnected systems, comprising numerous Electronic Control Units (ECUs) that communicate with each other. Ensuring secure communication between these components is essential to prevent unauthorised access and control.
  • Safety-Critical Systems: The integration of advanced driver-assistance systems (ADAS) and autonomous driving technologies means that any cybersecurity breach could directly impact vehicle safety. Ensuring the security of these systems is paramount to protecting the safety of passengers and other road users.
  • Physical Security: Automotive systems are subject to physical threats. Attackers can potentially gain physical access to vehicles, enabling them to tamper with hardware components or inject malicious code directly.
  • Over-the-Air (OTA) Updates: Many modern vehicles support OTA updates for software and firmware. While this feature enhances convenience and functionality, it also introduces new attack vectors.
  • User Privacy: Automotive systems collect and process a significant amount of data, including personal information, driving habits, and location data. Protecting user privacy is crucial to prevent misuse of this data.
In our previous work [7], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example. Our investigation yielded several findings, including:
@@
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security. However, STRIDE was not designed with automotive systems in mind, making its application in this context somewhat limited. Consequently, STRIDE often overlooks critical aspects of how cybersecurity breaches can directly affect automotive functions and safety. As a result, the reliability of the general outcomes from STRIDE is questionable when applied to the unique characteristics of automotive environments.
In other words, threat modelling of automotive systems often does not consider the consequences of cybersecurity attacks beyond information security. Similarly, attack models designed for traditional IT systems do not fully align with the multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions, necessitating a more comprehensive and dynamic approach to security modelling.
@@
Contributions. In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
Unlike existing security modelling approaches, ACTISM integrates various frameworks to dynamically update security models with new information, acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks. In order to bolster the completeness of the security model, ACTISM integrates the following threat analysis and risk assessment frameworks:
@@
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
@@
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities. CVSS (Common Vulnerability Scoring System): This model is also static as it provides a standardised method for evaluating the severity of vulnerabilities at a particular point in time. It does not dynamically adapt to new threats without manual intervention.
@@
Table 1 lists the differences between ACTISM and these security modelling frameworks, and their respective limitations. In contrast with other security modelling frameworks, the ACTISM framework integrates various frameworks to be dynamic, continuously adapting to new threats and updating the automotive system’s cybersecurity posture. This distinction sets it apart from the static models, providing a more robust approach to maintaining automotive cybersecurity.
Table 1: Comparison of ACTISM with other security modelling frameworks
Framework Type Limitations
ACTISM Dynamic Relies on the integration of various security modelling frameworks.
STRIDE Static Limited to specific threat categories. Does not dynamically adapt to new threats without manual intervention.
Attack Trees Static Requires manual updates and may not cover all emerging threats. Does not dynamically adapt to new threats without manual intervention.
HEAVENS Static Primarily focuses on high-level security analysis which may not cover detailed, system-specific vulnerabilities and threats. Does not dynamically adapt to new threats without manual intervention.
CVSS Static Does not dynamically adapt to new threats without manual intervention.
TARA Static Does not dynamically adapt to new threats without manual intervention.
Although frameworks such as STRIDE and CVSS were designed for IT security threats, they can also be applied to physical threats. As an example, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption. As an integrative approach, ACTISM provides guidance on how both cyber and physical attacks can be modeled using these frameworks to capture the various entry points and methods attackers might use, ensuring comprehensive security coverage across both cyber and physical domains.
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
Furthermore, we demonstrate the utility of ACTISM by illustrating its adaptive response to a real-world vulnerability disclosure affecting the IVI system. This case study underscored ACTISM’s effectiveness in addressing emerging threats and evolving security challenges in automotive cybersecurity. Furthermore, the study showcased the versatility of ACTISM, demonstrating its potential applicability to other automotive systems beyond the IVI, making it a robust tool for safeguarding a wide range of vehicle technologies.
Finally, we outline avenues for future research and development in this area, including automated vulnerability management workflows.
2 Related Work
Various works have attempted to address security modelling in automotive systems.
@@ 2 Related Work
Abuabed et al. [1] proposed a framework for assessing the cybersecurity vulnerabilities of automotive vehicles. Their approach employed STRIDE to classify threats. For assessing the impact of threats, they used version 2.0 of the HEAVENS security model [9], and they evaluated attack feasibility with CVSS version 2.0. Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS). Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths. This method facilitates a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
However, the framework has several limitations. Firstly, in the threat identification process, the authors did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This omission could be because of the assumption that physical attack vectors do not involve flows or exchanges of data. Notwithstanding this assumption, physical attack vectors can be exploited to facilitate data flows, such as "USB Man-in-the-Middle (USB MITM)" attacks, where an attacker intercepts and potentially alters the communication between a USB device and a host system. Consequently, physical attack vectors can be a crucial gap in the overall security design of the vehicle.
Secondly, the CVSS version 2.0 framework adopted in Abuabed et al. [1] is no longer current. The newer versions of CVSS include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [2] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. In addition, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [5] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
@@
3 Threat-informed Dynamic Security Modelling with ACTISM
@@
We propose ACTISM - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example and apply ACTISM to a real-world case study, where we demonstrate its ability to dynamically adapt to new threats and improve the security of automotive systems. This case study provides concrete evidence of the framework’s applicability and benefits, distinguishing our research from previous work.
Current security modelling frameworks offer valuable insights but lack adaptability to evolving cyber-physical threats. ACTISM addresses this gap by introducing a framework that continuously updates the cybersecurity posture of automotive systems in response to both cyber and physical threats, attack vectors, and system vulnerabilities.
@@
Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks. This iterative approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [15], breaching the trust boundary within security models.
@@
ACTISM does not replace current threat and attack modelling frameworks. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By integrating various security modelling frameworks into an iterative approach, ACTISM enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM comprises eight key steps as shown in Figure 1.
Figure 1: Overview of ACTISM
To enhance the practical utility of ACTISM among cybersecurity practitioners, we adopt the STRIDE methodology using a STRIDE-per-element approach, where threats are assessed against each element in the Data Flow Diagram. We choose not to use the STRIDE-per-interaction approach, as research has shown it yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [13]. Furthermore, we utilise HEAVENS version 2.0 [9] for impact assessment. To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155.
In addition, we employ CVSS version 3.1 to assess attack feasibility, which allows us to leverage contemporary cybersecurity vulnerability knowledge bases. CVSS version 3.1 provides granular metrics and scoring options, considering various factors that affect the exploitability and impact of a vulnerability. This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
@@
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
@@
4.1 Step 1: Identify assets
@@
In this experience report, we reference the IVI HPC system DFD from Das et al. (2024) [4] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling. It is important to note that this DFD is an example and may not encompass newer types of attacks, such as dirty road attacks that exploit the fact that autonomous vehicles rely on sensors and cameras to interpret their environment. Attackers may place objects or patterns on the road that confuse these sensors, leading to incorrect navigation decisions or even accidents.
@@
4.2 Step 2: Construct Data Flow Diagrams
@@
4.3 Step 3: Model threats
@@
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Table 9 [8].
4.4 Step 4: Evaluate impact
To evaluate the impact of the identified threats, we use the HEAVENS version 2.0 framework [9] which is specifically designed for automotive risk assessments.